# Self-Assembled Chains and Solids of Dipolar Atoms in a Multilayer

### v1

@@ Acknowledgements.
We gratefully acknowledge the fruitful discussions with Joaquim Casulleras. We acknowledge financial support from Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033 (Spain) under Grant No. PID2020-113565GB-C21 and from AGAUR-Generalitat de Catalunya Grant No. 2021-SGR-01411. GG and GM acknowledge partial support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), with the CRC-TRR 306 “QuCoLiMa”, Project-ID No. 429529648. GG acknowledges funding from the European Union, the Spanish Ministry of Universities and the Recovery, Transformation and Resilience Plan through a grant from Universitat Politècnica de Catalunya.
@@ II Diffusion Monte Carlo Method
The ground-state properties of the system are calculated using the diffusion Monte Carlo (DMC) method (for details of the method see, for example, Refs. Boronat and Casulleras (1994); Hammond et al. (1994); Toulouse et al. (2015)). This method has demonstrated its capacity to provide an accurate description of correlated quantum systems Ceperley (2010). Examples include ultradilute bosonic Cikojević et al. (2019); Parisi et al. (2019) and fermionic mixtures Comparin et al. (2019), few and many-bosonic bound-states Guijarro et al. (2020, 2022), Bose Ardila et al. (2020) and Fermi Bombín et al. (2019) polarons, dipolar Bose supersolid stripes Bombin et al. (2017), Bose gas subject to a multi-rod lattice Rodríguez-López et al. (2021), and ultracold quantum gases with spin-orbit interactions Sánchez-Baena et al. (2018). The DMC method solves the many-body Schrödinger equation in imaginary time, yielding a precise estimation of the ground-state energy within controllable statistical error. By using guiding wave functions, it is possible to impose a certain symmetry characteristic for a given phase and study phase transitions. Also, the use of guiding wave functions acts as in importance sampling, meaning that the more relevant regions of phase space are better sampled, minimizing the variance and the statistical error. To describe the gas phase we chose a translationally invariant trial wave function in Bijl-Jastrow (BJ) form, composed as a pair product of two-body correlation terms, while for the solid phase we used a Nosanow-Jastrow (NJ) wave function, incorporating one- and two-body correlations factors and breaking the translational symmetry. To describe the chain phase we introduce a new trial wave function, combining two-body correlation terms with many-body terms.
II.1 Trial wave function: Gas phase
@@
II.2 Trial Wave Function: Solid phase
@@
II.3 Trial wave function: Chain Gas phase
@@ II.4 Freezing density
II.5 Threshold density
The dipolar interlayer potential
$V_{\rm int}({r},l)=d^{2}\frac{\left(r^{2}-2l^{2}h^{2}\right)}{\left(r^{2}+l^{2}h^{2}\right)^{5/2}},\quad l=1,2,\ldots,M-1,$ (13)
is anisotropic. It is attractive for small values of $r$ but repulsive for large values of $r$, where $r$ is the in-plane distance between dipoles. Consider an $M$-layer system of dipolar bosons in a chain phase. Now, we ask ourselves what is the threshold distance $r_{\rm t}$ between a dipole on the first layer and a dipole on the last layer, and in a next chain such that the dipolar interlayer potential is zero. This threshold distance $r_{\rm t}$ can be obtained from Eq. (13)
$d^{2}\frac{r_{t}^{2}-2(M-1)^{2}h^{2}}{(r_{t}^{2}+(M-1)^{2}h^{2})^{5/2}}=0,$ (14)
with $l=M-1$. Now we solve for $r_{\rm t}$
$r_{\rm t}^{2}=2(M-1)^{2}h^{2}.$ (15)
In Fig. 5 we show the interlayer potential $V_{\rm int}(r,M-1)$ as a function of $r/r_{0}$ for different values of the number of layers $M$, with $h/r_{0}=1.0$. The threshold distances $r_{\rm t}$ are the points where the potential crosses the horizontal axis. The values of $r_{\rm t}$ are indicated by arrows in Fig. 5.
Figure 5: Interlayer potential $V_{\rm int}(r,M-1)$ as a function of $r/r_{0}$ for different values of the number of layers $M$, with $h/r_{0}=1.0$. The arrows show the positions of the threshold distances $r_{\rm t}$ for $M=$2, 3 and 4.
Now, the mean-inter-particle distance in one layer is proportional to
$\langle r/r_{0}\rangle\sim\frac{1}{\sqrt{nr_{0}^{2}/M}}\quad\to\quad\frac{nr_{0}^{2}}{M}\sim\frac{1}{\langle r/r_{0}\rangle^{2}}.$ (16)
From the Eq. (15) and Eq. (16) we obtain the threshold density $n_{\rm t}$, which satisfies Eq. (14)
$n_{\rm t}r_{0}^{2}=\frac{M}{2(M-1)^{2}(h/r_{0})^{2}}.$ (17)
The threshold density Eq. (17) is an approximation. Densities larger than the threshold density $n_{\rm t}r_{0}^{2}$ (attractive interlayer potential) will favor the formation of a gas phase and lower densities than the threshold density (repulsive interlayer potential) will favor the formation of the ordered phases, either solid or chains.
II.6 Finite size effects

### v2

@@ Acknowledgements.
We gratefully acknowledge the fruitful discussions with Joaquim Casulleras. We acknowledge financial support from Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033 (Spain) under Grant No. PID2020-113565GB-C21 and from AGAUR-Generalitat de Catalunya Grant No. 2021-SGR-01411. GG and GM acknowledge partial support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), with the CRC-TRR 306 “QuCoLiMa”, Project-ID No. 429529648. GG acknowledges funding from the European Union-Next Generation, the Spanish Ministry of Universities and the Recovery, Transformation and Resilience Plan through a grant from Universitat Politècnica de Catalunya.
End matter on trial wave functions. We use guiding wave functions to impose a specific symmetry on the state. We describe the gas phase by the Bijl-Jastrow (BJ) wave function,
$\Psi_{\rm G}(\mathbf{r}_{1},\dots,\mathbf{r}_{N})=\prod_{i<j}^{N}f_{2}(|\mathbf{r}_{i}-\mathbf{r}_{j}|),$ (2)
where $f_{2}(r)$ is a two-body correlation function. The solid is described by a Nosanow-Jastrow (NJ) wave function,
which contains one- and two-body terms $f_{1}(r)$ and $f_{2}(r)$, respectively. The key difference is that the BJ function maintains translational symmetry, while the NJ describes a state with a broken translational symmetry.
In addition to the gas and solid phases, the presence of anisotropic interactions allows for the emergence of a dipolar chain phase. This phase is similar to the molecular phase observed in bilayer systems Macia et al. (2014); Astrakharchik et al. (2016); Cinti et al. (2017); Guijarro et al. (2022) but extends beyond two layers. While the use of BJ (for gas) and NJ (for crystal) trial functions is well established in the literature, to the best of our knowledge, a Monte Carlo description for dipolar chain systems is still lacking. Here, we propose a new type of trial wave function specifically designed to describe chain phases. Its primary characteristic lies in not only capturing interlayer correlations but also enabling the free movement of chains within planes. In our model, a single chain consists of $M$ dipoles, each belonging to a different layer without any crystal ordering between them. The new trial wave function is constructed as follows:
where $f_{M}(|\mathbf{r}_{i}-\mathbf{R}_{i}|)$ is a many-body term, depending on the position of an atom $\mathbf{r}_{i}$ and the center of mass of the chain $\mathbf{R}_{i}=~{}(\sum_{k\in{\rm C}_{i}}\mathbf{r}_{k})/M$, with the sum covering all atoms belonging to the same chain as the $i$th atom. We consider $f_{M}(r)$ to have a Gaussian shape given by $f_{M}(r_{i})=e^{-\alpha|\mathbf{r}_{i}-\mathbf{R}_{i}|^{2}}$, where the variational parameter $\alpha$ quantifies the strength of localization. The function $f_{M}(r)$ is a many-body term, as a displacement of a single atom leads to $M$ changes in the function due to the resulting movement of the center of mass. A crucial difference between the solid and chain trial wave functions is that the positions of the lattice sites $\mathbf{R}_{i}$ in the NJ function remain static and are fixed by the triangular lattice, whereas they are allowed to change in $\Psi_{\rm C}$ during the simulation. We found that the chain trial wave function results in lower energy in certain regions of the phase diagram compared to the trial wave functions for solids and gases. Therefore, the novel introduction of the chain trial wave function is essential for fully describing the phase diagram of the dipolar multilayer system (further details can be found in the Supplemental Material Note (1)).
@@ II Diffusion Monte Carlo Method
The ground-state properties of the system are calculated using the diffusion Monte Carlo (DMC) method (for details of the method see, for example, Refs. Boronat and Casulleras (1994); Hammond et al. (1994); Toulouse et al. (2015)). This method has demonstrated its capacity to provide an accurate description of correlated quantum systems Ceperley (2010). Examples include ultradilute bosonic Cikojević et al. (2019); Parisi et al. (2019) and fermionic mixtures Comparin et al. (2019), few and many-bosonic bound-states Guijarro et al. (2020, 2022), Bose Ardila et al. (2020) and Fermi Bombín et al. (2019) polarons, dipolar Bose supersolid stripes Bombín et al. (2017), Bose gas subject to a multi-rod lattice Rodríguez-López et al. (2021), and ultracold quantum gases with spin-orbit interactions Sánchez-Baena et al. (2018). The DMC method solves the many-body Schrödinger equation in imaginary time, yielding a precise estimation of the ground-state energy within controllable statistical error. By using guiding wave functions, it is possible to impose a certain symmetry characteristic for a given phase and study phase transitions. Also, the use of guiding wave functions acts as in importance sampling, meaning that the more relevant regions of phase space are better sampled, minimizing the variance and the statistical error. To describe the gas phase we chose a translationally invariant trial wave function in Bijl-Jastrow (BJ) form, composed as a pair product of two-body correlation terms, while for the solid phase we used a Nosanow-Jastrow (NJ) wave function, incorporating one- and two-body correlations factors and breaking the translational symmetry. To describe the chain phase we introduce a new trial wave function, combining two-body correlation terms with many-body terms.
III Trial wave function: Gas phase
@@
IV Trial Wave Function: Solid phase
@@
V Trial wave function: Chain Gas phase
@@ VI Freezing density
VII Finite size effects